OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius. HELP PLEASE

OpenStudy (anonymous):

3000 = (h * l)+ pi * r^2 l = 2*pi*r 3000 = (h * 2*pi*r)+ pi * r^2

OpenStudy (anonymous):

3000 = (50 * 2*pi*r)+ pi * r^2 3000 = 50pir + pir^2 solve for r

OpenStudy (anonymous):

pi = 3.14 or 22/7

OpenStudy (anonymous):

ok I think i can figure it out. Thank you so much!

OpenStudy (anonymous):

HELP! I dont get it! What was the final answer for R?